Fermi National Laboratory


Accelerator Update for January 4, 2006 - January 6, 2006

Wednesday January 4

The day shift began with the Recycler Ring (RR) stash at 322.9 E10, the Antiproton Source (Pbar) stack at 49.4 E10. Beam is off to MiniBooNE, NuMI, and MCenter due to Tevatron (Tev) access for F17 Kicker and Collimator replacement.

At 8:37 AM the Main Injector (MI) pulsed power was turned off in order to investigate the feeder 86 fault.

Since no one is currently taking beam, some general maintenance is being done in Booster, MI, and MiniBoone.

At 9:59 AM the cause for the fault on feeder 86 was traced to an arcing transformer in Kautz Road utility yard.  The arc was caused by a mouse, repair work is beginning.

Transformer repair was completed at 3:04 PM. The MI is turning back on.

At 5:38 PM the F17 kicker and collimator replacement is complete.

At 6:15 PM a vacuum leak at F11 caused further delays in turning on Main Injector beam.  Access is being made into F sector to investigate.

The access into F sector was completed at 11:17 PM, a failed turbo pump was the cause of the poor vacuum.  Main Injector is being turned on.

F1 is now cooling down, and is expected to be cold by 7:00 AM.

Problems with a Booster power supply (B:MV10) were resolved at 11:54 PM.

Thursday January 5

The midnight shift began with the Recycler Ring (RR) stash at 303.3 E10, Pbar stack at 48.1 E10, and with stacking, MiniBooNE, NuMI, and MCenter turning back on. 

At 4:36 AM, an analyzing magnet power supply (F:MC7AN1) tripped off and caused MCenter downtime for the rest of the midnight shift.

At 5:37 AM PBar began experiencing rad trip, expert wre consulted and the problem was resolved by 8:00 AM.

At 8:36 AM, while trying to turn on the Tev, a ground fault was found in F-sector.

At 9:58 an access into F-Sector begins in search of the ground fault.  This access turns off stacking, NuMI, MiniBoooNE and MIPP.

The F-Sector ground fault problem was resolved by 4:00 PM. 

Beam returns to MiniBoooNE at 4:39 PM.

Beam to NuMI at 6:08 PM.

Stacking resumes at 7:18 PM, it was delayed by power supply problems.  MCenter is still off due to power supply problems with F:MC7AN1.  Experts are working on the supply.

Operations put the Recycler into shot setup at 8:35 PM to stash antiprotons.

At 9:55 PM the recycler stash established a new record of 370.1 E10.

Pbar returns to stacking at 10:11 PM

Friday January 6

The midnight shift began with the Tev in tune up mode, the Recycler Ring (RR) stash at 361.6 E10, and with MiniBooNE, and NuMI taking beam.

At 1201 AM, MCenter is turning on after repairs on F:MC7AN1 have been completed.

Shot setup begins at 1:34AM.

Operations established store 4581 at 3:34 AM with an initial luminosity of 171.85 E30.  This is a new world record.

Pbar returns to stacking at 4:52 AM, beam also returns to MCenter.

The Plan for Wednesday

Keep the store, stack, stash, run beam to NuMI, MiniBoone, and MCenter. 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/6/2006   email Fermilab